OpenStudy (anonymous):

how do you do y=(x-2)^2+3?? Please explain and help asap

OpenStudy (anonymous):

y = (x-2)^2 + 3 y = (x^2 - 4x + 4) + 3 y = x^2 -4x +7

OpenStudy (anonymous):

where do you get -4x?

OpenStudy (anonymous):

(x-2)^2 = (x-2)*(x-2) = x^2 - 2x - 2x + 4 = x^2 -4x + 4

OpenStudy (anonymous):

thank you!!!!!